OpenStudy (anonymous):

multiply. simplify answer as much as possible. (x-3)/(4x+12) * (x^2+x-6)/(x^2-4x+3)

OpenStudy (anonymous):

\[ \frac{(x-3) \left(x^2+x-6\right)}{(4 x+12) \left(x^2-4 x+3\right)} \] Is this your problem to be simplified?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ \frac{(x-3) \left(x^2+x-6\right)}{(4 x+12) \left(x^2-4 x+3\right)}=\frac{(x-3) (x-2) (x+3)}{4 (x+3) (x-3) (x-1)} \] Can you finish it now?

OpenStudy (anonymous):

no i still need help

OpenStudy (anonymous):

Cancel what is common up and down.

OpenStudy (anonymous):

\[ \frac{(x-3) \left(x^2+x-6\right)}{(4 x+12) \left(x^2-4 x+3\right)}=\frac{(x-3) (x-2) (x+3)}{4 (x+3) (x-3) (x-1)}=\frac{x-2}{4 (x-1)} \]

OpenStudy (anonymous):

Did you understand it?